OpenStudy (watermelon_talk):

Carrington is paid at the same rate each time he baby sits. The graph shows the rates he is paid for different times. Which equation represents the situation? A. B. C. D.

OpenStudy (watermelon_talk):

Carrington is paid at the same rate each time he baby sits. The graph shows the rates he is paid for different times. Which equation represents the situation? A. B. C. D.

OpenStudy (anonymous):

sorry, i kinda need options

OpenStudy (watermelon_talk):

A. Y=1/7X-5 B. Y=1/7X+5 C. Y=7X-5 D.Y=7X+5 @Pikachubacca

OpenStudy (anonymous):

is it math?